I think I'm starting to understand how to solve these types of equations. However, I think I may need a bit more "English" as to the exact steps. Could someone help me with solving x^-2 = 4.
Thanks!

Found 2 solutions by ankor@dixie-net.com, tutorcecilia:
Answer by ankor@dixie-net.com(12701)

solve x^-2 = 4.
Just follow the simple rules of algebra.
:
Negative exponent involves the reciprocal: x^-2 = 1/x^2
:
1/x^2 = 4
:
Get rid of x^2 in the denominator; mult equation by x^2 resulting in:
1 = 4x^2
:
Divide both sides by 4 resulting in:
1/4 = x^2
:
Or we can say: x = SqRt[1/4]
;
x = 1/2;

Answer by tutorcecilia(2144)

x^-2=4
The easiest way is just to flip a negative exponent:
x^-2=1/x^2 This will make the exponent positive without changing the value
So, 1/x^2=4
sqrt(1)/sqrt(x)^2=sqrt(4) [Take the square root of both sides]
1/x=2 [Solve for x]
(1/x)(x/1)=(x/1)(2)
1=2x
1/2=x
.
Check by plugging (x=1/2) back into the original equation:
x^-2=4
(1/2)^-2=4
(2/1)^2=4[flip to get a positive exponent]
4/1=4 [Simplify-
4=4 [checks out]